If I have 10x-7y=-42 how can I write it in a Slope-Intercept Form ?

Found 2 solutions by Alan3354, lwsshak3:
Answer by Alan3354(69443)

solve for y.

Answer by lwsshak3(11628)

If I have 10x-7y=-42 how can I write it in a Slope-Intercept Form ?
Slope-Intercept Form: y=mx+b, m=slope, b=y-intercept
10x-7y=-42
7y=10x+42
y=10x/7+6, slope=10/7, y-intercept=6s